OpenStudy (anonymous):

The height s of a ball (in feet) thrown with an initial velocity of 64 feet per second from an initial height of 7 feet is given as a function of the time t (in seconds) by s(t)=-16t^2+64t+7 Graph s(t) and answer; a)Determine the time at which height is maximum b)What is the maximum height.

OpenStudy (anonymous):

take derivative to this function that is  the t is the time

OpenStudy (anonymous):

I am suppose to be entering this in ti83 and finding the answer...

OpenStudy (anonymous):

The solutions of the quadratic equation ax2 + bx + c = 0 are x = when a ≠ 0 and b2 º 4ac ≥ 0. You can read this formula as “x equals the opposite of b, plus or minus the square root of b squared minus 4ac, all divided by 2a.” ºb ± b2º 4 ac 2a

OpenStudy (anonymous):

\[s^'(t)=0\] speed zero at max height

OpenStudy (anonymous):

\[s^'(t)=-(2)16t+64=0\]

OpenStudy (anonymous):

b)use the t above in \[s()=-16( )^2+64()+7\] empty space is yur time above

OpenStudy (anonymous):

Thats Right But You Look My Answer thats Really Right

OpenStudy (anonymous):

I am really confused, but here is an example. So is this close to what you guys are saying.

OpenStudy (anonymous):

its not confusing the question is basically asking for the Turning point (bvertex can you find that)

OpenStudy (anonymous):

did you do calculus derivatives and its application

OpenStudy (anonymous):

Jonask : Good Qouestion

OpenStudy (anonymous):

no, I don't think so

OpenStudy (anonymous):

ok so can you find a vertex of a parabola

OpenStudy (anonymous):

y=-16x^2+64x+7

OpenStudy (anonymous):

@Jonask so is it 2

OpenStudy (anonymous):

I think I figured this out. So I solved the quad. formula it came out to 2 and I plug the 2 in -16(2)^2+64(2)+7=71, which gives me the answer to the second part.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks for the help